OpenStudy (anonymous):

Multiply and simplify if possible. √7x (√x-7√7)

OpenStudy (anonymous):

thank you so much, and I really appreciate the way you are explaining the answer to me.

OpenStudy (solomonzelman):

\[\sqrt{7x} =\sqrt{x-7\sqrt{7}}\]THIS? OR what?

OpenStudy (solomonzelman):

Can you just draw it or use equation editor, because I don;t get the question.

OpenStudy (anonymous):

you got it correct, but there is no equal sign, and the ending is in para. strange isn't it. let me post it again

OpenStudy (solomonzelman):

\[\sqrt{7x}\sqrt{x-7\sqrt{7}}~~~like~~this?\]

OpenStudy (anonymous):

yes, that's it

OpenStudy (solomonzelman):

\[\sqrt{7x} \times \sqrt{x-7\sqrt{7}}\]I don;t think there is a way to simplify this.

OpenStudy (solomonzelman):

Actually, my bad....

OpenStudy (solomonzelman):

\[\sqrt{7x} \times \sqrt{x-7\sqrt{7}}~~~~->~~~~~\sqrt{\color{red} {{7x \times (x-7\sqrt{7})}}}\]

OpenStudy (anonymous):

ok, so it cannot be simplified, hmmm let me write this one.

OpenStudy (solomonzelman):

\[\sqrt{7x^2-49x \sqrt{7}}\]I think that's the most I can simplify, that's it, and that's the answer.

OpenStudy (solomonzelman):

It could be, see my replies?

OpenStudy (solomonzelman):

\[I~~USED:~~~\sqrt{a} \times \sqrt{b}=\sqrt{a \times b}\]

OpenStudy (anonymous):